OpenStudy (anonymous):

solve by factoring x^2-36=5x

hero (hero):

x^2 - 5x - 36 = 0 x^2 - 9x + 4x - 36 = 0 x(x - 9) + 4(x - 9) = 0 (x - 9)(x + 4) = 0 x - 9 = 0 x = 9 x + 4 = 0 x = -4

OpenStudy (anonymous):

thx